Twenty metrics of liquidity, Solvency, and Profitability The comparative financial statements of Automotive Solutions Inc. are...

  1. Twenty metrics of liquidity, Solvency, and Profitability

    The comparative financial statements of Automotive Solutions Inc. are as follows. The market price of Automotive Solutions Inc. common stock was $119.70 on December 31, 20Y8.

    AUTOMOTIVE SOLUTIONS INC.
    Comparative Income Statement
    For the Years Ended December 31, 20Y8 and 20Y7
        20Y8     20Y7
    Sales $10,000,000 $ 9,400,000
    Cost of goods sold (5,350,000) (4,950,000)
    Gross profit $ 4,650,000 $ 4,450,000
    Selling expenses $ (2,000,000) $(1,880,000)
    Administrative expenses (1,500,000) (1,410,000)
    Total operating expenses $ (3,500,000) $(3,290,000)
    Operating income $ 1,150,000 $ 1,160,000
    Other revenue and expense:
        Other revenue 150,000 140,000
        Other expense (interest) (170,000) (150,000)
    Income before income tax $ 1,130,000 $ 1,150,000
    Income tax expense (230,000) (225,000)
    Net income $ 900,000 $ 925,000
    AUTOMOTIVE SOLUTIONS INC.
    Comparative Statement of Stockholders’ Equity
    For the Years Ended December 31, 20Y8 and 20Y7
    20Y8 20Y7
    Preferred
    Stock
    Common
    Stock
    Retained
    Earnings
    Preferred
    Stock
    Common
    Stock
    Retained
    Earnings
    Balances, Jan. 1 $500,000 $500,000 $5,375,000 $500,000 $500,000 $4,545,000
    Net income 900,000 925,000
    Dividends:
        Preferred stock (45,000) (45,000)
        Common stock (50,000) (50,000)
    Balances, Dec. 31 $500,000 $500,000 $6,180,000 $500,000 $500,000 $5,375,000
    AUTOMOTIVE SOLUTIONS INC.
    Comparative Balance Sheet
    December 31, 20Y8 and 20Y7
        Dec. 31, 20Y8     Dec. 31, 20Y7
    Assets
    Current assets:
        Cash $ 500,000 $ 400,000
        Marketable securities 1,010,000 1,000,000
        Accounts receivable (net) 740,000 510,000
        Inventories 1,190,000 950,000
        Prepaid expenses 250,000 229,000
            Total current assets $3,690,000 $3,089,000
    Long-term investments 2,350,000 2,300,000
    Property, plant, and equipment (net) 3,740,000 3,366,000
    Total assets $9,780,000 $8,755,000
    Liabilities
    Current liabilities $ 900,000 $ 880,000
    Long-term liabilities:
        Mortgage note payable, 10% $ 200,000 $ 0
        Bonds payable, 10% 1,500,000 1,500,000
            Total long-term liabilities $1,700,000 $1,500,000
    Total liabilities $2,600,000 $2,380,000
    Stockholders' Equity
    Preferred $0.90 stock, $10 par $ 500,000 $ 500,000
    Common stock, $5 par 500,000 500,000
    Retained earnings 6,180,000 5,375,500
    Total stockholders' equity $7,180,000 $6,375,000
    Total liabilities and stockholders' equity $9,780,000 $8,755,000

    Instructions:

    Determine the following measures for 20Y8.

    Round ratio values to one decimal place and dollar amounts to the nearest cent. For number of days' sales in receivables and number of days' sales in inventory, round intermediate calculations to the nearest whole dollar and final amounts to one decimal place. Assume there are 365 days in the year.

    1. Working capital $
    2. Current ratio
    3. Quick ratio
    4. Accounts receivable turnover
    5. Days' sales in receivables days
    6. Inventory turnover
    7. Days' sales in inventory days
    8. Debt ratio %
    9. Ratio of liabilities to stockholders' equity
    10. Ratio of fixed assets to long-term liabilities
    11. Times interest earned times
    12. Times preferred dividends earned times
    13. Asset turnover
    14. Return on total assets %
    15. Return on stockholders’ equity %
    16. Return on common stockholders’ equity %
    17. Earnings per share on common stock $
    18. Price-earnings ratio
    19. Dividends per share of common stock $
    20. Dividend yield %
Answer #1
Metrics Result Details Formula
1. Working capital $2,790,000 =3690000-900000 Current Assets Less Current Liabilities
2. Current ratio                       4.10 =3690000/900000 Current Assets Divided by Current Liabilities
3. Quick ratio                       2.50 =(500000+1010000+740000)/900000 Cash + Marketable Securities + Accounts Receivable/ Current Liabilities
4. Accounts receivable turnover                     16.00 =10000000/((740000+510000)/2) Sales / ((Opening Accounts Receivable + Closing Accounts Receivable) / 2)
5. Days' sales in receivables                     27.01 days =740000/10000000*365 Accounts Receivables / Sales * 365
6. Inventory turnover                       9.35 =10000000/((1190000+950000)/2) Sales / ((Opening Inventory + Closing Inventory) / 2)
7. Days' sales in inventory                     43.44 days =1190000/10000000*365 Inventory / Sales * 365
8. Debt ratio 19% % =1700000/(1700000+7180000) Debt / (Debt + Equity)
9. Ratio of liabilities to stockholders' equity                       0.36 =2600000/7180000 Total Liabilities / Total Stockholders Equity
10. Ratio of fixed assets to long-term liabilities                       1.44 =3740000/2600000 Total Fixed Assets / Total Long term liabilities
11. Times interest earned                       7.65 times =(1150000+150000)/170000 (Operating Income + Other Revenues) / Interest cost
12. Times preferred dividends earned                     20.00 times =900000/45000 Net Income / Preferred Dividend
13. Asset turnover                       2.81 =10000000/((3740000+3366000)/2) Sales / ((Opening Fixed Assets + Closing Fixed Assets) / 2)
14. Return on total assets 9.7% % =900000/((9780000+8755000)/2) Net Income / ((Opening Total Assets + Closing Total Assets) / 2)
15. Return on stockholders’ equity 13.3% % =900000/((7180000+6375000)/2)
16. Return on common stockholders’ equity 13.6% % =(900000-45000)/(((7180000-500000)+(6375000-500000))/2) Net Income less preferred dividend / total number of shares / (Opening total equity less preferred stock+closing total equity less preferred stock)/2
17. Earnings per share on common stock                       8.55 =(900000-45000)/100000 Net Income less preferred dividend / total number of shares. Total number of shares calculated as paid up share capital / par value per share (500000/5)
18. Price-earnings ratio                     14.00 =119.7/8.55 Share price as on 31 Dec 2018 / earning per share
19. Dividends per share of common stock                       0.50 =(50000)/100000 Total Dividend / total common shareholders equity
20. Dividend yield 3.57% % =0.5/14 Dividend per share / EPS